OpenStudy (anonymous):

Identify the amplitude and the period of the function c(x) = 3cos (x + π).

OpenStudy (mertsj):

Amplitude = 3, period = 2 pi

OpenStudy (phi):

if we call the period T, amplitude A, and phase P match your equation to \[ A \cdot cos( \frac{2 \pi}{T} x +P) \] A is easy, it is 3 to find T, match 2pi/T to the coefficient of x \[ \frac{2 \pi}{T} = 1 \] solve for T, to get T= \( 2\pi \) and, though not asked, the phase is \( \pi \)